MnO2 + 4HCl → MnCl2 + Cl2 + 2H2O
If 39.1 grams of MnO2 reacts with 48.2 g of HCl, and 19.8 grams of Cl2 was collected, what is the percent yield?
1
Expert's answer
2020-12-14T14:40:09-0500

First the limiting reactant should be determined:


"n(MnO_2)=\frac{39.1g}{86.94g/mol}=0.450 mol"


"n(HCl)=\frac{48.2g}{36.46g/mol}=1.32 mol"


"\frac{1.32}{0.450}=2.93<4" , hence HCl is the limiting reactant (1 mole of MnO2 requires 4 moles of HCl for the reaction). Therefore, theoretical yield of Cl2 will be:


"m(Cl_2)=1.32molHCl\times\frac{1molCl_2}{4molHCl}\times\frac{70.9gCl_2}{1molCl_2}=23.4g"


"\%yield=\frac{19.8g}{23.4g}\times100=84.6\%"


Answer: 84.6%